# Cosmic Ray Superdiffusion and Mirror Diffusion in Partially Ionized and Turbulent Medium

### v1

@@ 2.3 CRs’ perpendicular superdiffusion induced by wandering magnetic field lines
(2) When $\langle\lambda_{\parallel}\rangle<L_{\rm inj}$: the scattering of CRs along magnetic field lines is non-negligible, and their propagation becomes diffusive rather than ballistic, as numerically tested in (Hu et al., 2022). Parallel diffusion can be influenced predominantly by either pitch-angle scattering diffusion or bouncing from magnetic mirrors.
@@
Magnetic mirrors created by changes of magnetic field strength were assumed to trap CR with sufficiently large pitch angles, preventing them from further propagation unless their pitch angle decreases through resonant scattering Cesarsky & Kulsrud (1973). However, a recent study in Lazarian & Xu (2021) suggested that in MHD turbulence, the perpendicular superdiffusion of turbulent magnetic field lines prevents CRs from becoming trapped between magnetic mirrors, as the return of CR to the initial mirror is improbable. Instead, CRs repeatedly bounce among various magnetic mirrors while diffusing along the local magnetic field lines, leading to a new type of diffusion mechanism termed ”mirror diffusion” (Lazarian & Xu, 2021). According to Lazarian & Xu (2021), the transition between the scattering and mirror diffusion regimes occurs at a critical cosine of pitch angle $\mu_{c}$. For CRs with a pitch angle cosine $\mu$ greater than $\mu_{c}$, parallel diffusion is governed primarily by resonant pitch angle scattering, resulting in scattering diffusion; conversely, for $\mu$ less than $\mu_{c}$, mirror diffusion prevails. The mirror interactions of CRs are strong as they reverse the CRs’ motion. Thus, mirror diffusion is typically slower than that induced by resonance scattering Lazarian & Xu (2021). An example of the scattering diffusion and the mirror diffusion is given in Fig. 1.
A particle with a Larmor radius of 10 cells and an initial pitch angle cosine of 0.25 was injected into simulation A4 (with $M_{s}=1.24$ and $M_{A}=1.21$). Initially, the particle entered a region dominated by a magnetic field mirror, which caused it to oscillate back and forth. Over time, the particle gradually diffuses out of this mirror, accompanied by changes in its pitch angle. Eventually, the particle became dominated by scattering diffusion and entered another magnetic mirror region. Both scattering and mirror diffusion contribute to the reduction in the CRs’ diffusion perpendicular to the mean magnetic field. In this context, the perpendicular diffusion can be expressed as follows (Lazarian & Yan, 2014; Hu et al., 2022):
$\langle z^{2}\rangle\sim\frac{D_{\|}t^{3/2}}{27L_{\rm inj}}M_{A}^{4},~{}~{}{M_{A}\leq 1}$ (6)
$\langle z^{2}\rangle\sim\frac{D_{\|}t^{3/2}}{27L_{\rm inj}}M_{A}^{3},~{}~{}{M_{A}>1}$
@@
where $D_{\parallel}$ is the parallel diffusion coefficient. The transport of CRs is thus highly sensitive to the value of their mean free path $\langle\lambda_{\parallel}\rangle$, which is influenced by magnetic field fluctuations, for instance via pitch angle scattering.
@@
Strongly coupled regime: In the strongly coupled regime, ions and neutrals develop MHD turbulence due to sufficient collision and energy transfer. Suppose magnetic field fluctuations are significant in the strongly coupled regime such that $\langle\lambda_{\parallel}\rangle<L_{\rm inj}$, CRs experience strong scattering and propagate in a superdiffusive manner. Their perpendicular diffusion follows the scaling described by Eq. 5, with $\sqrt{\langle z^{2}\rangle}\propto t^{3/4}$.
@@
Intermediate coupling and fully decoupled regimes: When transitioning into the intermediately coupled or fully decoupled regimes, the damping of MHD turbulence occurs and magnetic field fluctuations are reduced, leading to an increase in $\langle\lambda_{\parallel}\rangle$. As a result, CRs shift into the $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$ regime. In this regime, CRs experience insignificant scattering, and their superdiffusion shifts to the scaling described by Eq. 4, with $\sqrt{\langle z^{2}\rangle}\propto t^{3/2}$, which is driven by the Richardson diffusion of magnetic field lines. The transition between these regimes highlights the critical role of ions and neutrals coupling in determining CR propagation.
@@
Figure 2: Left three panels: $x-z$ slices of ion velocity ($v_{i}$; top) and neutral velocity ($v_{n}$; bottom) at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the kinetic energy spectrum. The theoretically expected parallel and perpendicular components of the neutral-ion decoupling wavenumber are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, for comparison with Kolmogorov and Burgers turbulence scaling, respectively.
@@
Figure 3: Left three panels: $x-z$ slices of ion density ($\rho_{i}$; top) and neutral density ($\rho_{n}$; bottom) at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the density spectrum. The neutral-ion decoupling wavenumber’s theoretically expected parallel and perpendicular components are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, for comparison with Kolmogorov and Burgers turbulence scaling, respectively.
@@ 3.1 Two-fluid simulations
The scale-free 3D two-fluid simulations analyzed in this work were performed with the AthenaK code (Stone et al., 2024). We model a two-fluid magneto-fluid system consisting of an ionized fluid (ions plus electrons) and a neutral fluid (Hu et al., 2024). The simulation domain is discretized on a 12003 uniformly staggered grid. Under the assumptions of an isothermal equation of state with constant sound speed $c_{s}$ and periodic boundary conditions, we solve the ideal two-fluid MHD equations:
@@
where $\rho$ and $\mathbf{v}$ denote the mass density and velocity of the ionized (subscript ”$i$”) and neutral (subscript ”$n$”) fluids, respectively, and $\mathbf{B}$ is the magnetic field. The drag coefficient $\gamma_{\rm d}$ couples ions and neutrals through the collisional exchange of momentum, while the stochastic forcing terms $\mathbf{f}_{i}$ and $\mathbf{f}_{n}$ drive turbulence in both fluids with weights proportional to their densities.
@@
Figure 4: Left three panels: $x-z$ slices of magnetic field strength at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the magnetic field spectrum. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, respectively.
Figure 5: The parallel mean free path $\lambda_{\parallel}$ in the unit of turbulence injection scale $L_{\rm inj}$ as a function of $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, which is the ratio of the neutral-ion collisional frequency to the eddy turnover rate.
@@ 4.1.1 Velocity distribution and kinetic energy spectrum
Strongly coupled regime: Fig. 2 presents 2D slices of the velocity fields and turbulent kinetic energy spectra for ions and neutrals, taken perpendicular to the mean magnetic field at the center of the simulation box. In both transonic and supersonic conditions, the velocity structures of ions (upper panel) and neutrals (lower panel) are highly similar in the strongly coupled regime, closely resembling the single-fluid approximation. For the transsonic case, the ion and neutral velocity spectra approximately follow the Kolmogorov scaling, exhibiting a spectral slope of -5/3. In contrast, the supersonic case yields steeper spectra with slopes close to -2, indicative of Burgers scaling. These steeper spectra result from energy dissipation associated with shocks
@@
Intermediate coupling: In the intermediate coupling regime, the 2D velocity slices and spectra of ions and neutrals remain similar to each other, but both spectra become steeper than the Kolmogorov and Burgers scalings. This steepening indicates significant damping of the turbulent cascade. In this regime, neutrals begin to decouple from ions, although the decoupling is incomplete due to the fluctuations in ionization fractions. Consequently, neutrals cannot fully establish an independent hydrodynamic cascade, resulting in persistently steep spectra  [^4] The reduced value of $\gamma_{\rm d}$ and a high ionization fraction of 0.1 may cause enhanced ion inertial and frictional damping. However, a low ionization fraction 0.01 test has been given in the Appendix in Hu et al. (2024), showing that the neutral kinetic energy spectrum remains steep in the intermediate coupling regime.. This behavior suggests that neutral-ion decoupling occurs gradually over a range of scales rather than sharply at a specific scale. Compared to the strongly coupled regime, the velocity distributions of both neutrals and ions exhibit a clear deficiency of small-scale structures.
@@
In addition to the reason that fluctuations in ionization fraction result in different local decoupling scale, this regime is also likely to have a contribution from the anisotropy of MHD turbulence, which means Alfvén waves propagate predominantly along the magnetic field, while the turbulent cascade develops preferentially in the perpendicular direction (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999). As a result, neutrals may decouple from ions along the parallel direction yet remain coupled in the perpendicular direction due to turbulent motion. This behavior implies that neutral-ion coupling has both parallel and perpendicular components. Consequently, the decoupling scale is anisotropic, with its parallel component potentially much larger than the perpendicular component (Xu et al., 2015; Hu et al., 2024). In contrast, the classical model of neutral-ion coupling (see Eq. 1) considers only Alfvén waves and thereby omits the perpendicular coupling.
@@
Decoupled regime: In the decoupled regime, the ion and neutral spectra remain steep up to a wavenumber of approximately 30, which is close to the perpendicular decoupling scale. Beyond this scale, the neutral spectrum becomes shallower, indicating that neutrals have fully decoupled from ions and developed an independent hydrodynamic turbulent cascade characterized by a Kolmogorov slope. This trend was not observed in the earlier study by Hu et al. (2024) due to limited numerical resolution. Meanwhile, ions experience frequent collisions with neutrals, effectively damping ion turbulence. We find that the damping is more severe in the transonic case with the even steeper spectra slope than that in the supersonic case. It is likely that shock compression locally increases the density fluctuations and ionization fraction, efficiently reducing the decoupling. Although the velocity slices of ions and neutrals exhibit noticeable differences, both species predominantly retain only large-scale fluctuations.
@@ 4.1.2 Density distribution and density spectrum
Fig. 3 presents 2D density slices and corresponding density spectra for ions and neutrals. Under supersonic conditions, the density structures of ions (upper panel) and neutrals (lower panel) remain broadly similar across all coupling regimes, although subtle differences emerge. As decoupling progresses, small-scale density structures become less pronounced. In the strongly coupled regime, the neutral density spectra closely resemble those of ions, with both spectra exhibiting slopes shallower than the Kolmogorov scaling due to shock-induced small-scale density fluctuations. As the coupling weakens, the spectral slopes gradually steepen, particularly for neutrals, reflecting increased damping of small-scale density fluctuations.
@@
In contrast, under transonic conditions, the density spectra exhibit distinct characteristics. In the strongly coupled regime, both ion and neutral density spectra generally follow the Kolmogorov scaling, similar to their velocity spectra. However, as decoupling occurs, the ion density structures diverge significantly from those of neutrals, despite the continued similarity in their velocity fields. Ion density distributions exhibit pronounced small-scale striations, which are absent in the neutral density distributions. Correspondingly, the ion density spectrum remains shallow, whereas the neutral density spectrum steepens noticeably. This difference likely arises because the velocity field is dominated by incompressible Alfvénic turbulence, while compressible turbulent motions primarily drive density fluctuations. Although neutrals remain coupled to Alfvénic turbulent motions, their coupling to compressible MHD turbulence weakens, preventing the formation of small-scale density structures induced by compressible motions. In the fully decoupled regime, filamentary structures in the ion density vanish, and the ion density spectrum steepens significantly beyond a wavenumber of approximately 10. The observed differences between supersonic and transonic conditions are most likely attributable to shocks, as shock compression enhances density fluctuations and increases the ionization fraction, thereby strengthening ion-neutral coupling.
@@
Figure 6: Plots of perpendicular displacement of the particles $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}/L_{\rm inj}$ versus the CRs’ gyro periods $t\cdot\Omega$ in transonic $M_{s}\approx 1$ conditions. $l_{\rm diss}$ is the numerical dissipation scale and $L_{\rm inj}$ is the injection scale. The panels (a), (b), and (c) correspond to initial cosine of pitch angle $\mu\approx 1$, $\mu\approx 0.25$, and $\mu\approx 0.1$, respectively.

### v2

@@ 2.3 CRs’ perpendicular superdiffusion induced by perpendicular superdiffusion of magnetic field lines
(2) When $\lambda_{\parallel}<L_{\rm inj}$: the pitch-angle scattering and mirroring of CRs along magnetic field lines are non-negligible, and their propagation becomes diffusive rather than ballistic, as numerically tested in (Hu et al., 2022; Zhang & Xu, 2023).
@@
Magnetic mirrors created by changes of magnetic field strength were assumed to trap CR with sufficiently large pitch angles, preventing them from further propagation unless their pitch angle decreases through resonant scattering Cesarsky & Kulsrud (1973). However, a recent study in Lazarian & Xu (2021) suggested that in MHD turbulence, the perpendicular superdiffusion of turbulent magnetic field lines prevents CRs from being trapped between magnetic mirrors, as the return of CR to the initial mirror along the same field line is improbable, due to the field line stochasticity. Instead, CRs bounce among different magnetic mirrors while undergoing superdiffusion perpendicular to the mean field, leading to a diffusive motion along the magnetic field. This new parallel diffusion mechanism is termed ”mirror diffusion” (Lazarian & Xu, 2021). According to Lazarian & Xu (2021), the transition between the scattering and mirror diffusion regimes occurs at a critical cosine of pitch angle $\mu_{c}$, corresponding to the balance between the scattering and mirroring rates. For CRs with a pitch angle cosine $\mu$ greater than $\mu_{c}$, parallel diffusion is governed primarily by gyroresonant pitch angle scattering, resulting in scattering diffusion; conversely, for $\mu$ less than $\mu_{c}$, mirror diffusion prevails. The mirror interactions of CRs are strong, as they reverse the CRs’ motion. Thus, mirror diffusion is typically slower than that induced by resonance scattering Lazarian & Xu (2021). An example of the scattering diffusion and the mirror diffusion is given in Fig. 1.
A particle with a Larmor radius of 10 cells and an initial pitch angle cosine of 0.25 is injected into simulation A4 (with $M_{s}=1.24$ and $M_{A}=1.21$). Initially, the particle enters the mirroring-dominated diffusion regime due to its large pitch angle, which causes it to bounce back and forth as it interacts with multiple mirrors. Over time, the particle gradually decreases its pitch angle due to scattering. Then, the particle parallel transport becomes dominated by scattering diffusion and later again undergoes the mirror diffusion due to the stochastic change of pitch angle. In the mirror diffusion regions, the cosine of pitch angle is quasiperiodic, while in the scattering diffusion regions, the cosine of pitch angle is close to constant. Similar effects have been seen in Lemoine (2023). Both scattering and mirror diffusion can contribute to the CR parallel diffusion behavior, and thus further affect the perpendicular superdiffusion behavior. If $\lambda_{\parallel}<L_{\rm inj}$, the CR traveled along the magnetic field line by a distance of $x^{2}\sim D_{\parallel}t$, the thus perpendicular superdiffusion can be expressed as follows (Lazarian & Yan, 2014; Hu et al., 2022; Lazarian et al., 2023):
$\langle z^{2}\rangle\sim\frac{(D_{\|}t)^{3/2}}{27L_{\rm inj}}M_{A}^{4},~~{M_{A}\leq 1}$ (6)
$\langle z^{2}\rangle\sim\frac{(D_{\|}t)^{3/2}}{27L_{\rm inj}}M_{A}^{3},~~{M_{A}>1}$
@@
where $D_{\parallel}$ is the parallel diffusion coefficient. The transport of CRs is thus highly sensitive to the value of their mean free path $\lambda_{\parallel}$, which is influenced by magnetic field fluctuations.
@@
Strongly coupled regime: In the strongly coupled regime, ions and neutrals together carry MHD turbulence through sufficient collisions and energy transfer. When CRs’ Larmor radius is larger than the maximum parallel decoupling scale, they experience more significant magnetic field fluctuations than in the other decoupling regimes. The large fluctuations and thus higher scattering efficiency can cause a fast transition between scattering diffusion and mirror diffusion. Their perpendicular superdiffusion follows the scaling described by Eq. 5, with $\sqrt{\langle z^{2}\rangle}\propto t^{3/4}$.
@@
Intermediate coupling and fully decoupled regimes: In the intermediate or fully decoupled regimes, the damping of MHD turbulence occurs, and magnetic field fluctuations are reduced, leading to an increase in $\lambda_{\parallel}$. In this regime, CRs experience insignificant scattering, and the superdiffusive behavior of magnetic field lines may be altered, at scales smaller than the maximum parallel decoupling scales. If the initial perpendicular separation of CRs is smaller than the maximum parallel decoupling scale, their separation will grow more slowly before entering the perpendicular superdiffusion regime described by Eq. 4, with $\sqrt{\langle z^{2}\rangle}\propto t^{3/2}$. This superdiffusion is driven by the Richardson diffusion of magnetic field lines. The transition between these regimes underscores the critical role of ion–neutral coupling in regulating CR transport behavior.
In addition, the damping of MHD turbulence also erases small-scale magnetic field mirrors. CRs can be captured by large-scale magnetic field mirrors. The large spatial separation within the mirrors could be sufficient for CRs entering the perpendicular superdiffusion regime and undergoing mirror diffusion.
@@
Figure 2: Left three panels: xzx-z slices of ion velocity ($v_{i}$; top) and neutral velocity ($v_{n}$; bottom) at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the kinetic energy spectrum. The theoretically expected parallel and perpendicular components of the neutral-ion decoupling wavenumber are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, for comparison with Kolmogorov and Burgers turbulence scaling, respectively.
@@
Figure 3: Left three panels: xzx-z slices of ion density ($\rho_{i}$; top) and neutral density ($\rho_{n}$; bottom) at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the density spectrum. The neutral-ion decoupling wavenumber’s theoretically expected parallel and perpendicular components are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, for comparison with Kolmogorov and Burgers turbulence scaling, respectively.
@@ 3.1 Two-fluid simulations
The scale-free 3D two-fluid simulations analyzed in this work are performed with the AthenaK code (Stone et al., 2024). We model a two-fluid magneto-fluid system consisting of an ionized fluid (ions plus electrons) and a neutral fluid (Hu et al., 2024). The simulation domain is discretized on a 12003 uniformly staggered grid. Under the assumptions of an isothermal equation of state with constant sound speed $c_{s}$ and periodic boundary conditions, we solve the two-fluid MHD equations:
@@
where $\rho$ and $\mathbf{v}$ denote the mass density and velocity of the ionized (subscript ”ii”) and neutral (subscript ”nn”) fluids, respectively, and $\mathbf{B}$ is the magnetic field. $\gamma_{\rm d}$ is the drag coefficient, while the stochastic forcing terms $\mathbf{f}_{i}$ and $\mathbf{f}_{n}$ drive turbulence in both fluids with weights proportional to their densities.
@@
Figure 4: Left three panels: xzx-z slices of magnetic field strength at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the magnetic field spectrum. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, respectively.
Figure 5: The parallel mean free path $\langle\lambda_{\parallel}\rangle$ in the unit of turbulence injection scale $L_{\rm inj}$ as a function of coupling strength $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, which is the ratio of the neutral-ion collisional frequency to the eddy turnover rate at injection scale.
@@ 4.1.1 Velocity distribution and kinetic energy spectrum
Strongly coupled regime: Fig. 2, corresponding to the simulations A1, A3, A4, B1, B3, and B4, presents 2D slices of the velocity fields and turbulent kinetic energy spectra for ions and neutrals. The slices are taken perpendicular to the mean magnetic field at the center of the simulation box. In both transonic and supersonic conditions, the velocity structures of ions (upper panel) and neutrals (lower panel) are highly similar in the strongly coupled regime, closely resembling the single-fluid approximation. For the transsonic case, the ion and neutral velocity spectra approximately follow the Kolmogorov scaling, exhibiting a spectral slope of -5/3. In contrast, the supersonic case yields steeper spectra with slopes close to -2, indicative of Burgers scaling. These steeper spectra result from energy dissipation associated with shocks.
@@
Intermediate coupling: In the intermediate coupling regime, the 2D velocity slices and spectra of ions and neutrals remain similar to each other, but both spectra become steeper than the Kolmogorov and Burgers scalings. This steepening indicates significant damping of the turbulent fluctuations. In this regime, neutrals begin to decouple from ions in some local regions, although the decoupling is incomplete due to the fluctuations in ionization fractions. Consequently, neutrals cannot fully establish an independent hydrodynamic cascade, resulting in persistently steep spectra. The reduced value of $\gamma_{\rm d}$ and a high ionization fraction of 0.1 may cause enhanced ion inertial and frictional damping. However, a low ionization fraction 0.01 test is given in the Appendix in Hu et al. (2024), showing that the neutral kinetic energy spectrum remains steep in the intermediate coupling regime. This behavior suggests that neutral-ion decoupling occurs gradually over a range of scales rather than sharply at a specific scale due to the spatial variation of ionization fraction. Compared to the strongly coupled regime, the velocity distributions of both neutrals and ions exhibit a clear deficiency of small-scale structures.
@@
In addition to the reason that fluctuations in ionization fraction result in different local decoupling scales, this regime is also likely to have a contribution from the anisotropy of MHD turbulence, which means Alfvén waves propagate predominantly along the magnetic field, while the turbulent cascade develops preferentially in the perpendicular direction (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999). As a result, neutrals may decouple from ions along the parallel direction yet remain coupled in the perpendicular direction due to turbulent motion. This behavior implies that neutral-ion coupling has both parallel and perpendicular components. Consequently, the decoupling scale is anisotropic, with its parallel component potentially much larger than the perpendicular component (Xu et al., 2015; Hu et al., 2024). In contrast, the classical model of neutral-ion coupling (see Eq. 1) considers only Alfvén waves and thereby omits the perpendicular coupling.
@@
Decoupled regime: In the decoupled regime, the ion and neutral spectra remain steep up to a wavenumber of approximately 30, which is close to the perpendicular decoupling scale. Beyond this scale, the neutral spectrum becomes shallower, indicating that neutrals have fully decoupled from ions and developed an independent hydrodynamic turbulent cascade characterized by a Kolmogorov slope. This trend was not observed in the earlier study by Hu et al. (2024) due to limited numerical resolution. Meanwhile, ions still experience frequent collisions with neutrals, effectively damping MHD turbulence in ions. We find that the damping is more severe in the transonic case with an even steeper spectra slope than that in the supersonic case. Likely, compressible slow and fast modes contribute differently in transonic and supersonic conditions, or shock locally alters ionization fraction, efficiently reducing the decoupling. Although the velocity slices of ions and neutrals exhibit noticeable differences, both species predominantly retain only large-scale fluctuations.
@@ 4.1.2 Density distribution and density spectrum
Fig. 3 presents 2D density slices and density spectra for ions and neutrals, for the simulations A1, A3, A4, B1, B3, and B4. Under supersonic conditions, the density structures of ions (upper panel) and neutrals (lower panel) remain broadly similar across all coupling regimes, although subtle differences emerge. As decoupling progresses, small-scale density structures become less pronounced. These density structures are more likely associated with turbulent cascade, which falls in different coupling regimes. In the strongly coupled regime, the neutral density spectra closely resemble those of ions, with both spectra exhibiting slopes shallower than the Kolmogorov scaling due to shock-enhanced small-scale density fluctuations. As the coupling weakens, the spectral slopes gradually steepen, but only slightly, as the shock compression does not suffer from damping. This steepening is more apparent for neutrals, reflecting increased damping of small-scale density fluctuations.
@@
In contrast, under transonic conditions, where density fluctuations are primarily generated by the turbulent cascade, the density spectra exhibit distinct characteristics. In the strongly coupled regime, both ion and neutral density spectra generally follow a Kolmogorov-like scaling, consistent with their velocity spectra. However, as decoupling sets in, the ion density structures diverge markedly from those of the neutrals, despite the continued similarity of their velocity fields. The ion density distribution develops pronounced small-scale striations that are absent in the neutral counterpart. Correspondingly, the ion density spectrum remains close to the Kolmogorov scaling, whereas the neutral density spectrum steepens substantially.
This difference likely arises because the velocity field is dominated by incompressible Alfvénic turbulence, while density fluctuations are primarily driven by compressible turbulent motions. Although neutrals remain coupled to Alfvénic modes in the intermediate regime, their coupling to compressible MHD turbulence weakens, suppressing the formation of small-scale density structures associated with compressive motions. It is also possible that, when neutrals decouple from ions, the plasma $\beta$ of the ion fluid decreases. In this case, the sudden excess of magnetic fluctuation energy in the ion component can be converted into turbulent kinetic energy, further enhancing ion density fluctuations. While, over the relevant length scales, both magnetic and turbulent energy spectra steepen due to damping, the density fluctuations generated by the change in plasma $\beta$ can persist, allowing the ion density spectrum to retain a Kolmogorov-like form even in the intermediate and weakly coupled regimes. In the fully decoupled regime, filamentary structures in the ion density vanish, and the ion density spectrum steepens sharply beyond a wavenumber of approximately 10. The differences between the supersonic and transonic cases are most likely attributable to shocks: in supersonic turbulence, shock compression enhances density contrasts and thereby strengthens ion–neutral coupling.
@@
Figure 6: Plots of perpendicular displacement of the particles $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}/L_{\rm inj}$ versus the CRs’ gyro periods $t\cdot\Omega$ in transonic $M_{s}\approx 1$ conditions, corresponding to the simulations A1, A3, and A4. $l_{\rm diss}$ is the numerical dissipation scale and $L_{\rm inj}$ is the injection scale. The panels (a), (b), and (c) correspond to initial cosine of pitch angle $\mu\approx 1$, $\mu\approx 0.25$, and $\mu\approx 0.1$, respectively.